# Template AASArticle with Examples: v6.3.1111Released on March, 1st, 2021

### v1

@@ 5 Conclusions
Here we have described a technique to generate non-LTE observables from a multidimensional SN Ia model, motivated by the fact that spherically symmetric models and LTE calculations are insufficient to produce observables that match observation to a precise degree. We show that steady-state 1D calculations reproduce, to good approximation, the results from a time-dependent 2D calculation along corresponding lines of sight. As a result, this makes it possible to produce reasonably approximate multidimensional non-LTE spectra using 1D non-LTE calculations along individual lines of sight. An important component of this strategy is the constraint of the outgoing bolometric luminosity in each of the 1D calculations to match that of the 2D LTE calculation at corresponding lines of sight, which is necessary primarily due to the varying amounts of radioactive material in the ejecta profiles for different lines of sight. This technique is important as non-LTE calculations are essential for the calculation of a number of SNe Ia features and relations, including the $B$-band Phillips relation, but are often prohibitively expensive if not reduced in accuracy in multiple dimensions. In the double detonation model, for example, the asymmetry of the ejecta is an essential aspect as it allows for a variety of observable indicators based on the viewing angle to a given explosion, including peak magnitude and expansion velocity, which also vary among the population of observed SNe Ia. This technique allows for high-accuracy non-LTE observables from relatively efficient calculations while preserving the asymmetry that is inherent to the model, improving the unification between theory and observation.
@@
While the observables highlighted in this work show general improvement, issues still remain for this model, particularly at northern lines of sight. This includes Si ii $\lambda$6355 velocities and B-V color that are somewhat higher than observed, in addition to slightly small $\Delta M_{B,15}$ values. It is possible that these issues may be alleviated by 3D modeling of the double detonation scenario as the polar wedges from our 2D models are most affected by the imposed cylindrical symmetry. Additionally, effects from the companion star, which are not considered in the model examined in this paper, can alter the amount of diversity across lines of sight (Boos et al., 2024; Pollin et al., 2024).
Future work on this technique will involve a closer examination of the origin of the multidimensional deviations of the observables (e.g. Figure 2) to determine the quality of approximation of this strategy. We will also use this technique on a wider range of models. This includes double detonation models with masses above and below 1.0 M, in addition to double detonation models that involve a companion. Non-LTE observables from models that include an exploding companion (i.e. triple or quadruple detonation; Papish et al. 2015; Tanikawa et al. 2019; Pakmor et al. 2022; Pollin et al. 2024) will also be particularly interesting due to recent work that has shown, in LTE, that events arising from one or two star explosions in this scenario can have remarkably similar spectra (Boos et al., 2024). Other sub-Chandrasekhar SN Ia models may also be used with this technique to determine how they fare in comparison to the population of observed SNe Ia.
We thank Stefan Taubenberger for his reduction of the Hicken et al. (2009) data. S.J.B. and D.M.T. acknowledge support from the National Science Foundation under Grant No. 2307442. S.J.B. also acknowledges support from NASA grant HST AR-16156. K.J.S. is supported by NASA through the Astrophysics Theory Program (80NSSC20K0544). This work was granted access to the HPC resources of TGCC under the allocation 2023 – A0150410554 on Irene-Rome made by GENCI, France. Resources supporting this work were also provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center and the University of Alabama high performance computing facility.
Appendix A Spectra
@@
We show the 1D LTE spectra for all wedges at 15 and 30 days post-explosion, compared with the 2D LTE and 1D non-LTE spectra in Figures 13 and 14, respectively.
Appendix B Data
We provide the 2D ejecta and 15 wedges profiles associated with the our examined model on Zenodo (10.5281/zenodo.14004460). Also included in this repository is our synthetic observables from each of our calculations (2D LTE, 1D LTE, and 1D non-LTE), as well as our pseudo-2D non-LTE observables. This includes spectra and UBVRI light curves between -3 and +15 d from $B$-band peak, with a cadence of 1.5 d.

### v2

@@ 5 Conclusions
Here we have described a technique to generate non-LTE observables from a multidimensional SN Ia model, motivated by the fact that spherically symmetric models and LTE calculations are insufficient to produce observables that match observation to a precise degree. We show that 1D steady-state calculations reproduce, to good approximation, the results from a time-dependent 2D calculation along corresponding lines of sight. As a result, this makes it possible to produce reasonably approximate multidimensional non-LTE spectra using 1D non-LTE calculations along individual lines of sight. An important component of this strategy is the constraint of the outgoing bolometric luminosity in each of the 1D calculations to match that of the 2D LTE calculation at corresponding lines of sight, which is necessary primarily due to the varying amounts of radioactive material in the ejecta profiles for different lines of sight. This technique is important as non-LTE calculations are essential for the calculation of a number of SNe Ia features and relations, including the $B$-band Phillips relation, but are often prohibitively expensive if not reduced in accuracy in multiple dimensions. In the double detonation model, for example, the asymmetry of the ejecta is an essential aspect as it allows for a variety of observable indicators based on the viewing angle to a given explosion, including peak magnitude and expansion velocity, which also vary among the population of observed SNe Ia. This technique allows for high-accuracy non-LTE observables from relatively efficient calculations while preserving the asymmetry that is inherent to the model, improving the unification between theory and observation.
@@
While the observables highlighted in this work show general improvement, issues still remain for this model, particularly at northern lines of sight. This includes Si ii $\lambda$6355 velocities and B-V color that are somewhat higher than observed, in addition to slightly small $\Delta M_{B,15}$ values. It is possible that these issues may be alleviated by 3D modeling of the double detonation scenario as the polar wedges from our 2D models are most affected by the imposed cylindrical symmetry. Additionally, effects from the companion star, which are not considered in the model examined in this paper, can alter the amount of diversity across lines of sight (S. J. Boos et al., 2024; J. M. Pollin et al., 2024).
Future work on this technique will involve a closer examination of the origin of the multidimensional deviations of the observables (e.g. Figure 2) to determine the quality of approximation of this strategy. Additionally, it would be fruitful to establish avenues to reduce the uncertainty brought upon by using different codes between the non-LTE and LTE calculations. The disparity between the LTE (Sedona) and non-LTE (CMFGEN) results in this work are qualitatively similar to what is shown by S. Blondin et al. (2022), which reviewed the consistency between several radiative transfer codes for SNe Ia. As different radiative transfer codes can produce varied results, using a standardized method that can isolate specific physical processes, while still considering multidimensional effects, would improve the reliability and interpretation of the results generated from our technique.
We will also use this technique on a wider range of models. This includes double detonation models with masses above and below 1.0 M, in addition to double detonation models that involve a companion. Non-LTE observables from models that include an exploding companion (i.e. triple or quadruple detonation; O. Papish et al. 2015; A. Tanikawa et al. 2019; R. Pakmor et al. 2022; J. M. Pollin et al. 2024) will also be particularly interesting due to recent work that has shown, in LTE, that events arising from one or two star explosions in this scenario can have remarkably similar spectra (R. Pakmor et al., 2022; S. J. Boos et al., 2024; J. M. Pollin et al., 2024). Other sub-Chandrasekhar SN Ia models may also be used with this technique to determine how they fare in comparison to the population of observed SNe Ia.
We thank the referee for their helpful comments. We also thank Stefan Taubenberger for his reduction of the M. Hicken et al. (2009) data. S.J.B. and D.M.T. acknowledge support from the National Science Foundation under Grant No. 2307442. S.J.B. also acknowledges support from NASA grant HST AR-16156. K.J.S. is supported by NASA through the Astrophysics Theory Program (80NSSC20K0544). This work was granted access to the HPC resources of TGCC under the allocation 2023 – A0150410554 on Irene-Rome made by GENCI, France. Resources supporting this work were also provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center and the University of Alabama high performance computing facility.
Appendix A Full Spectral Sets
@@
We show the 1D LTE spectra for all wedges at 15 and 30 days post-explosion, compared with the 2D LTE and 1D non-LTE spectra in Figures 15 and 16, respectively.
Appendix B NIR Spectra
Here we highlight the IR spectra from our calculations, in comparison with SN 2011fe. The observed spectrum shown is corrected for distance, redshift, and Galactic reddening. The 2D LTE and pseudo-2D non-LTE synthetic spectra are rebinned beyond 10,000 Å due to the poor Monte-Carlo sampling at these wavelengths.
Figure 17: Optical-NIR spectrum of SN 2011fe (E. Y. Hsiao et al., 2013) compared with our 1D non-LTE and 2D LTE spectra for wedge 12 of our model. The cyan shading indicates the difference between 1D non-LTE and pseudo-2D non-LTE spectra at the given time. Spectra are labeled based on the time relative to their respective $B$-band maximum.
We show an NIR spectral comparison between SN 2011fe and that from wedge 12 of our calculations in Figure 17. Around maximum light, the 1D non-LTE spectra provides a satisfactory match with SN 2011fe in the NIR out to $\sim 13000$, which includes a number of Ca ii and Mg ii features (see E. Y. Hsiao et al. 2013 for a detailed discussion of the NIR spectral features of SN 2011fe). However, the modest pseudo-2D non-LTE deviation from the 1D non-LTE suggests that multidimensional effects may be important in the NIR. Beyond $13000$, none of our model spectra are able to well-replicate the flux or features observed in SN 2011fe at max light, or 3.7 days post-max. By 8.2 days, however, generally improved agreement can be observed in most of the NIR region displayed between SN 2011fe and the 1D non-LTE spectra. The improved agreement between SN 2011fe and the 1D non-LTE calculation is more stark at 12.3 days, but the pseudo-2D non-LTE is notably deviated from the 2D LTE beyond $15000$.
Appendix C Alternative Power Injection Technique
An alternative method to increase the outgoing luminosity of the 1D LTE calculations to match that of the 2D LTE treatment was used in a previous version of this work, which we describe here. In this method, rather than scaling the radioactive decay energies, we introduce an artificial power source at the origin of the grid. The luminosity of this source is set such that the total outgoing luminosity of a given calculation in 1D matches the 2D results of the corresponding line of sight. The source is a blackbody, with the results being sensitive to the choice of temperature ($3\times 10^{4}$ K is used here). Additional testing showed that the power source could be placed out to upwards of $\sim 5,000$ km s-1 without significant alterations to the spectra.
Figure 18: 1D spectra at all valid lines of sight at 15 and 30 days after the explosion. Spectra from our standard set of 1D LTE calculations, where the decay energy is scaled, are shown in black. The orange shading indicates how spectra under the alternative method, where a power source is placed at the origin, deviates from our standard strategy. The lines of sight are labeled in orange, corresponding with the wedges in Figure 1. The three southernmost lines of sight are not shown at 30 days post-explosion as their luminosities are unable to be unified.
The spectra of the two 1D LTE methods are compared in Figure 18. At 15 days post-explosion, the two strategies produce nearly identical spectra. By 30 days post-explosion, the spectra are still broadly very similar for most wedges, except for the region around 5500 Å. Our decay energy scaling method is more effective in reproducing the 2D spectra in this region at this time (see Figure 15). Spectra at 30 days post-explosion for wedges 13-15 are not shown in Figure 18 as these spectra have an unmodified bolometric luminosity in steady-state that is higher than the target 2D result, with no functional way to reduce the outgoing luminosity in the power source strategy.
The difference in physical motivation and effectiveness between these two techniques is fairly marginal. However, the decay energy scaling strategy has an advantage over the other because of its ability to reduce the outgoing luminosity in a straightforward way (i.e. a decay energy scaling factor below unity). This is necessary for the unification of bolometric luminosities with the 2D calculation for the three southernmost wedges at later times, which cannot be obtained through the power source method as it can only increase the outgoing luminosity.
Appendix D Data
We provide the 2D ejecta and 15 wedges profiles associated with the our examined model on Zenodo (10.5281/zenodo.15122976). Also included in this repository is our synthetic observables from each of our calculations (2D LTE, 1D LTE, and 1D non-LTE), as well as our pseudo-2D non-LTE observables. This includes spectra and UBVRI light curves between -3 and +15 d from $B$-band peak, with a cadence of 1.5 d.